OpenStudy (anonymous):

solve by completing the square. a^2-6a+_____

OpenStudy (dtan5457):

value of c= (b/2)^2

OpenStudy (dtan5457):

your value of b here is -6

OpenStudy (dtan5457):

(-6/2)^2=________?

OpenStudy (anonymous):

my options are 6 3 -3 and 9

OpenStudy (dtan5457):

ok. -3^2 is what?

OpenStudy (danjs):

(-3)^2

OpenStudy (anonymous):

9?

OpenStudy (danjs):

-3^2 = -9

OpenStudy (dtan5457):

yes and my bad. it is (-3)^2

OpenStudy (danjs):

(-3)^2 = 9

OpenStudy (dtan5457):

thank you @DanJS

OpenStudy (anonymous):

thank you guys